If (u) = 11, (v) = 23, and (u-v) = 30.

Find (u+v)


1
Expert's answer
2022-05-02T15:22:14-0400

The Law of Cosines


"30^2=11^2+23^2-2(11)(23)\cos \theta"

"\cos \theta=-\dfrac{125}{253}"

"|\vec{u}+\vec{v}|^2=11^2+23^2+2(11)(23)\cos \theta"

"|\vec{u}+\vec{v}|=\sqrt{11^2+23^2-2(11)(23)(\dfrac{125}{253})}"

"|\vec{u}+\vec{v}|=20"